Circular No. 6604

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1737–31 W. Cui and E. H. Morgan, Massachusetts Institute of Technology;

W. A. Heindl, University of California, San Diego; J. H. Swank, Goddard Space Flight Center; and D. M. Smith, University of California, Berkeley, report: “Preliminary results of the RXTE observation noted on IAUC 6603 reveal remarkable similarities to the blackhole candidate Cyg X-1 in its hard (or low) state. The power-density spectrum is flat below ∼ 0.03 Hz and roughly inversely proportional to the frequency above, with an integrated rms variability of 21.4 percent. The energy spectrum can be characterized by a power law with photon index ∼ 1.6. A more detailed analysis is required to determine if an iron line at ∼ 6.4 keV is accounted for by the diffuse x-ray emission in the central region of the galactic plane. The bestfit hydrogen column density is 4–6 × 10^{22} cm^{2}, depending on the diffuse background, in any case consistent with a location near the galactic center. Fluxes in the bands 2–10, 10–35, 35–75, and 75–150 keV were 18, 37, 60, and 90 mCrab, respectively. Assuming a distance of 8.5 kpc, the 2–150keV source luminosity was 1.9 × 10^{37} erg/s. Similarities to Cyg X-1 make GRS 1737–31 a likely black hole candidate. Further observations at all wavelengths, in particular, infrared and radio, are encouraged.”

COMET C/1995 O1 (HALE-BOPP) N. Dello Russo, M. A. DiSanti, and M. J. Mumma, Goddard Space

Flight Center, NASA; K. Magee-Sauer, Rowan College of New Jersey; R. Novak, Iona College; and M. Fomenkova, University of California at San Diego, report: “We detected water in C/1995 O1 on Jan. 21.5, Feb. 23.9– 24.1, and Mar. 1.7–1.9 UT, using the NASA Infrared Telescope Facility (+ CSHELL cryogenic infrared spectrometer) on Mauna Kea. For Feb. 23.9, line fluxes (× 10^{−17} W m^{−2}, in an aperture 1′′ × 1′′.4 centered on the nucleus) were as follows: 111–100 band and 202–303 line, 1.2; 100–010 band and 101-110 line, 7.2. Five lines were analyzed. Lines of the 001– 010 vibrational band were also detected but show the expected effect of an optically-thick pump and so were not used in these reductions. The waterproduction rate (× 10^{30} molecules/s) was 2.7 on Jan. 21.5, 4.3 on Feb. 23.9, and 4.0 on Mar. 1.8, assuming a rotational temperature of 70 K and an ortho-para-ratio of 2.45. After detailed data reduction, we report that spectral lines of HDO (100 vibrational band) were not detected, contrary to the report on IAUC 6573. The preliminary reductions were adversely affected by high and variable telluric HDO.”

1997 March 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green